A discrete random variable has probability mass function, P(X = n) = 12 n.
Let Y =  1, for x even
 1, for x odd
Find the expected value of Y ; (E[y]).
1
Expert's answer
2020-06-02T20:16:56-0400

Py(1)=P(x is even) 1/2n even when n multiples of 4, so P(x is even)=1/4=0.25 - every fourth number

Py(-1)=1-0.25=0.75

E(Y)=0.25*1+0.75*(-1)=-0.50